OpenStudy (anonymous):

Solve this system of linear equations. Separate the x- and y-values with a comma. -7x = -94 -8y 5x = -70 – 8y

OpenStudy (anonymous):

-7x = -94 - 8y 5x = -70 = 8y rearrange

OpenStudy (anonymous):

-7x = -94 - 8y on rearranging becomes -7x + 8y = -94

OpenStudy (anonymous):

5x = -70 - 8y on rearranging becomes 5x + 8y = -70

OpenStudy (anonymous):

so now we have -7x + 8y = -94 5x + 8y = -70

OpenStudy (anonymous):

Nevermind I don't need help bye

syed98 (syedmohammed98):



syed98 (syedmohammed98):

Bye

OpenStudy (anonymous):

BYE!